# Explainable Benchmarking for Iterative Optimization Heuristics

### 31 Jan 2024 v1

@@ 4.3.2. Algorithm/configuration selection
30 24 True (0.00) sobol (0.01) False (0.00) False (0.03) 20.0 (0.03) IPOP (0.00) mirrored (0.00) 10.0 (0.01) psr (0.01) default (0.01) 0.384299 -
@@
The exact settings for each single-best, as well as the avg-best configuration, are shown in Table 5. This table highlights that, while the contribution trend of a module might be negative in Figure 2 or 3, it might still be beneficial in combination with other modules, or its impact might be sufficiently low when performance is high that changing its setting does not change the performance significantly. This can be observed for example with the ‘covariance’ parameter, which as mentioned before seems to have a positive contribution to performance when turned on. However, for 21 out of 50 settings, the best-performing configuration is an MA-ES instead of CMA-ES variant. Since Table 5 also shows the impact of each module choice relative to the average of the other choices, we can isolate the performance gained from enabling or disabling the covariance adaptation in more detail. We observe that in the cases where this adaptation is turned off, the normalized AOCC would be lowered by between $0.00$ and $0.02$ in 16 out of 21 cases, while the reverse change would lead to much larger deterioration in performance.
@@
Other patterns, such as the impact of mirrored sampling, are more consistent between the two types of analysis. Table 5 shows that all of the best configurations enable this module, with a slight preference for ‘mirrored’ over ‘mirrored pairwise’. The changes in performance from the single-module changes also highlight the low impact of restart strategies on unimodal problems. From looking only at the selected modules, we might infer that ‘BIPOP’ is needed to achieve optimal performance on F1, but since restarts are not triggered in this scenario, the performance of our selected configuration would not change if restart is changed to ‘IPOP’, or even turned off completely.
@@ 4.4.1. Contribution of hyperparameters
Figure 4. Hyper-parameter contributions to the AOCC (shap values) per benchmark function for d=5 for modular DE. Categorical parameters are integer encoded in the visualisation in alphabetical order. For example, ‘mutation_base’ is either ‘best’ (0), ‘rand’ (1) or ‘target’ (2). Boolean parameters such as ‘lpsr’ are shown with low feature value for false (0) and high for true (1).
Figure 5. Hyper-parameter contributions per benchmark function for d=30 for modular DE. Categorical parameters are integer encoded in the visualisation in alphabetical order. For example, ‘mutation_base’ is either ‘best’ (0), ‘rand’ (1) or ‘target’ (2). Boolean parameters such as ‘lpsr’ are shown with low feature value for false (0) and high for true (1)
For the modular Differential Evolution, we have a total of $36288$ configurations. We should note that, similar to the CMA-ES, the models from which we calculate the SHAP-values are intentionally overfit, resulting in $R^{2}$ scores between $0.9$ and $0.96$. Figures 4 and 5 show these resulting SHAP values for all functions, in dimensionality 5 and 30 respectively. For dimensionality 5, we observe strong effects from the population size parameter $\lambda$, with the default (smallest) value often contributing negatively to the anytime performance. For the unimodal problems, the population size $2\cdot d$ seems to be preferred, while the multimodal functions benefit from an increase to $10\cdot d$. For the adaptation method, we surprisingly find that turning this module off generally leads to good performance, while performance deteriorates when jDE is enabled.
@@
When looking at the discretized parameters F and CR, we observe some interesting effects. While for most functions, the impact changes seemingly in trend with the size of the parameter, e.g., F becoming more positive the lower the value becomes, some 30-dimensional functions show completely different types of impact for CR. As an example, F7 shows that the highest CR value has both a very positive and a very negative impact on the anytime performance. This highlights an inherent limitation with the choice of SHAP as our analysis method, in that we only focus on the single-module impacts. In this version of DE, is seems intuitive that several modules might interact with each other, such as the choice of crossover method and CR, which might explain the unusual SHAP plot for F7. Within IOH-Xplainer we can zoom in on these kinds of higher-order interactions by utilizing different explainability methods, we consider this outside the scope of this current paper.
@@ 4.4.2. Algorithm/configuration selection
Another interesting observation from Table 8 is that the highest population sizes are never chosen in the 30-dimensional setting. This is likely influenced by the combination of chosen budget (10 0001000010\,00010 000) and performance measure, where very high population sizes could lead to slower, but eventually more precise, convergence. This highlights that, as with any benchmarking setup, the design choices made should be carefully considered when drawing conclusions from IOH-Xplainer.
@@
Table 7. Performance summary of modular DE. The gain avg-best is the average improvement of the average best configuration over all configurations. The gain single-best is the average improvement of the single-best configurations (per function) over all configurations. Boldface in the table indicates a significant improvement (avg-best versus all, and single-best versus avg-best).
@@
Table 8. Hall of fame of modular DE: best configurations per function including average best (denoted by All in fid), with resulting values of AOCC and structural BIAS indicator. Denoted in brackets after each module setting is the average effect in AOCC score of choosing the respective option versus all other choices (keeping the rest of the settings the same).
@@
5 1 0.25 (0.25) 0.25 (0.29) nan (0.12) bin (0.01) 8 (0.06) False (0.19) best (0.34) 1 (0.00) rand (0.02) True (0.55) 0.971233 gaps/clusters
5 2 0.25 (0.37) 0.25 (0.45) nan (0.14) bin (0.01) 8 (0.13) False (0.38) best (0.31) 2 (0.00) rand (0.01) True (0.11) 0.940875 bounds
5 3 0.05 (0.40) 0.5 (0.59) nan (0.13) exp (0.00) 8 (0.33) False (0.59) target (0.10) 2 (0.18) rand (0.08) True (0.06) 0.885237 -
@@
5 7 1.0 (0.35) 0.5 (0.44) nan (0.08) bin (0.09) 50 (0.78) False (0.04) target (0.18) 1 (0.24) pbest (0.33) False (0.17) 0.909025 gaps/clusters
@@
5 9 0.75 (0.53) 0.5 (0.51) nan (0.06) exp (0.01) 10 (0.43) True (0.01) target (0.29) 1 (0.09) pbest (0.16) True (0.66) 0.799625 -
5 10 1.0 (0.60) 0.25 (0.56) nan (0.30) exp (0.04) 10 (0.49) True (0.01) target (0.76) 2 (0.25) nan (0.74) True (0.39) 0.777871 -
5 11 1.0 (0.57) 0.25 (0.54) nan (0.27) exp (0.01) 10 (0.42) True (0.01) target (0.76) 2 (0.17) nan (0.69) True (0.26) 0.832987 -
@@
5 13 1.0 (0.49) 0.75 (0.55) nan (0.20) bin (0.00) 50 (0.44) True (0.08) target (0.47) 1 (0.33) best (0.24) False (0.51) 0.714799 bounds
@@
5 16 1.0 (0.30) 0.25 (0.32) nan (0.25) exp (0.08) 8 (0.22) True (0.29) best (0.33) 2 (0.10) rand (0.14) True (0.09) 0.466607 bounds
5 17 1.0 (0.33) 0.5 (0.34) nan (0.16) exp (0.03) 50 (0.31) True (0.03) target (0.24) 1 (0.29) pbest (0.25) False (0.26) 0.713114 centre
5 18 1.0 (0.33) 0.5 (0.35) nan (0.19) bin (0.01) 50 (0.31) True (0.02) target (0.25) 1 (0.30) pbest (0.29) False (0.25) 0.664711 centre
5 19 0.75 (0.04) 0.25 (0.05) jDE (0.06) exp (0.05) 10 (0.05) False (0.03) best (0.05) 1 (0.04) nan (0.05) True (0.09) 0.259360 bounds
5 20 0.25 (0.25) 0.5 (0.26) shade (0.22) exp (0.13) 10 (0.37) True (0.16) rand (0.29) 1 (0.11) nan (0.21) True (0.37) 0.678659 bounds
@@
5 23 0.75 (0.04) 0.25 (0.05) shade (0.05) exp (0.04) 50 (0.06) True (0.04) best (0.05) 2 (0.03) nan (0.05) False (0.04) 0.236681 bounds
5 24 1.0 (0.03) 0.5 (0.05) nan (0.02) bin (0.01) 50 (0.03) True (0.00) target (0.04) 1 (0.05) best (0.02) False (0.04) 0.123116 gaps/clusters
@@
30 2 0.25 (0.22) 0.75 (0.38) nan (0.19) bin (0.19) 14 (0.38) False (0.32) target (0.32) 1 (0.24) pbest (0.13) False (0.14) 0.565194 -
30 3 0.05 (0.02) 0.5 (0.01) shade (0.01) exp (0.02) 14 (0.09) False (0.08) rand (0.03) 1 (0.02) nan (0.03) True (0.00) 0.422156 -
30 4 0.05 (0.02) 0.5 (0.01) shade (0.01) exp (0.01) 14 (0.08) False (0.07) rand (0.03) 1 (0.02) nan (0.02) True (0.00) 0.405671 -
30 5 0.25 (0.32) 1.75 (0.03) nan (0.29) bin (0.05) 14 (0.23) False (0.11) best (0.02) 2 (0.14) best (0.02) True (0.03) 0.970231 bounds
@@
30 7 0.75 (0.03) 0.5 (0.06) nan (0.01) bin (0.04) 60 (0.03) True (0.01) target (0.06) 1 (0.04) pbest (0.03) True (0.01) 0.416182 -
30 8 0.25 (0.06) 0.5 (0.12) nan (0.01) bin (0.04) 14 (0.13) False (0.10) target (0.07) 2 (0.01) pbest (0.03) True (0.02) 0.375220 -
@@
30 11 1.0 (0.04) 0.25 (0.05) nan (0.02) exp (0.00) 14 (0.01) True (0.03) target (0.03) 2 (0.01) pbest (0.02) True (0.01) 0.392524 centre
@@
30 16 0.75 (0.02) 0.25 (0.03) nan (0.01) bin (0.02) 14 (0.02) False (0.02) best (0.03) 2 (0.01) nan (0.00) True (0.01) 0.433995 bounds
30 17 0.5 (0.03) 0.5 (0.05) nan (0.01) bin (0.05) 60 (0.02) True (0.01) target (0.04) 1 (0.04) pbest (0.03) False (0.00) 0.497815 -
30 18 0.75 (0.03) 0.5 (0.04) nan (0.01) bin (0.04) 60 (0.01) True (0.01) target (0.04) 1 (0.03) pbest (0.03) False (0.01) 0.459689 centre
@@
30 20 0.25 (0.10) 0.25 (0.21) nan (0.02) bin (0.02) 14 (0.14) False (0.08) best (0.02) 2 (0.01) rand (0.02) False (0.00) 0.472318 -
@@
30 23 1.0 (0.01) 0.25 (0.01) shade (0.02) exp (0.01) 14 (0.02) False (0.02) best (0.02) 2 (0.01) nan (0.01) True (0.03) 0.481032 bounds

### 23 Feb 2024 v2

@@ 4.3.2. Algorithm/configuration selection
5 2 true (0.03) Sobol (0.02) false (0.01) false (0.91) 8 (0.20) IPOP (0.00) mirrored (0.04) 5 (0.00) CSA (0.03) $\lambda$-decay (0.46) 0.907190 centre
5 3 true (0.11) Sobol (0.18) true (0.19) false (0.16) 20 (0.27) BIPOP (0.15) pairwise (0.21) 5 (0.30) CSA (0.19) equal (0.19) 0.378857 centre
5 4 false (0.00) Gaussian (0.04) true (0.06) false (0.05) 10 (0.05) BIPOP (0.04) mirrored (0.04) 10 (0.04) CSA (0.08) $\lambda$-decay (0.10) 0.168084 centre
5 5 true (0.00) Gaussian (0.00) true (0.01) false (0.00) 5 (0.02) IPOP (0.00) mirrored (0.00) 5 (0.00) PSR (0.00) default (0.03) 0.998051 centre
5 6 false (0.00) Halton (0.01) true (0.92) true (0.05) 5 (0.00) BIPOP (0.00) pairwise (0.00) 2 (0.02) CSA (0.11) $\lambda$-decay (0.01) 0.952707 centre
5 7 false (0.09) Halton (0.04) true (0.71) true (0.85) 20 (0.14) BIPOP (0.00) pairwise (0.05) 10 (0.06) CSA (0.04) equal (0.16) 0.951310 centre
5 8 true (0.02) Sobol (0.06) true (0.01) false (0.06) 8 (0.20) IPOP (0.00) mirrored (0.04) 5 (0.01) CSA (0.07) $\lambda$-decay (0.27) 0.909659 centre
5 9 false (0.02) Sobol (0.05) false (0.06) false (0.84) 5 (0.00) IPOP (0.01) mirrored (0.04) 2 (0.72) CSA (0.11) default (0.06) 0.915234 centre
5 10 false (0.00) Sobol (0.20) true (0.04) true (0.90) 5 (0.30) IPOP (0.00) mirrored (0.02) 5 (0.07) PSR (0.04) equal (0.47) 0.903269 centre
5 11 true (0.10) Sobol (0.38) true (0.02) false (0.63) 20 (0.71) IPOP (0.00) mirrored (0.10) 10 (0.47) CSA (0.02) default (0.05) 0.917868 centre
5 12 true (0.03) Halton (0.05) false (0.08) false (0.85) 8 (0.23) BIPOP (0.00) mirrored (0.07) 5 (0.01) CSA (0.06) $\lambda$-decay (0.37) 0.859747 centre
5 13 true (0.00) Sobol (0.22) true (0.00) true (0.90) 5 (0.76) IPOP (0.00) mirrored (0.77) 5 (0.64) PSR (0.15) equal (0.47) 0.895403 centre
5 14 false (0.30) Sobol (0.01) true (0.77) true (0.78) 5 (0.13) IPOP (0.00) pairwise (0.01) 5 (0.02) CSA (0.01) equal (0.03) 0.954450 centre
5 15 false (0.00) Halton (0.13) true (0.08) false (0.32) 200 (0.00) BIPOP (0.00) mirrored (0.13) 100 (0.29) PSR (0.05) default (0.30) 0.481335 centre
5 16 true (0.18) Sobol (0.11) true (0.37) false (0.38) 20 (0.43) BIPOP (0.08) mirrored (0.12) 5 (0.43) CSA (0.31) default (0.22) 0.809630 centre
5 17 false (0.03) Sobol (0.02) true (0.02) false (0.10) 20 (0.45) BIPOP (0.00) mirrored (0.05) 10 (0.19) CSA (0.08) default (0.11) 0.897028 centre
5 18 true (0.08) Sobol (0.15) true (0.21) false (0.37) 20 (0.56) BIPOP (0.15) mirrored (0.23) 5 (0.31) PSR (0.14) default (0.31) 0.818370 centre
5 19 true (0.13) Halton (0.11) false (0.12) false (0.14) 20 (0.18) IPOP (0.00) pairwise (0.10) 10 (0.12) CSA (0.12) default (0.12) 0.402468 centre
5 20 true (0.06) Sobol (0.10) true (0.11) false (0.16) 200 (0.19) BIPOP (0.00) mirrored (0.06) 20 (0.14) PSR (0.03) default (0.15) 0.356620 centre
5 21 false (0.47) Halton (0.50) true (0.21) true (0.34) 10 (0.34) IPOP (0.02) pairwise (0.25) 5 (0.37) PSR (0.19) default (0.19) 0.844145 centre
5 22 false (0.58) Halton (0.34) true (0.43) true (0.21) 5 (0.28) BIPOP (0.03) pairwise (0.12) 5 (0.17) PSR (0.03) $\lambda$-decay (0.28) 0.793936 bounds
5 23 true (0.22) Halton (0.14) true (0.18) false (0.48) 20 (0.42) BIPOP (0.02) mirrored (0.23) 5 (0.42) PSR (0.22) default (0.21) 0.664302 centre
5 24 false (0.01) Sobol (0.01) true (0.03) false (0.04) 10 (0.03) IPOP (0.01) mirrored (0.03) 5 (0.04) CSA (0.03) default (0.02) 0.141390 centre
30 All true (0.01) Sobol (0.05) false (0.00) false (0.16) 14 (0.00) IPOP (0.01) mirrored (0.02) 7 (0.01) CSA (0.01) default (0.01) 0.496598 centre
30 1 false (0.00) Sobol (0.17) false (0.01) false (0.06) 5 (0.18) IPOP (0.00) mirrored (0.02) 5 (0.01) PSR (0.01) $\lambda$-decay (0.34) 0.945324 centre
30 2 true (0.00) Sobol (0.07) false (0.01) true (0.22) 10 (0.08) IPOP (0.00) pairwise (0.00) 10 (0.04) PSR (0.21) equal (0.10) 0.294408 gaps/clusters
30 3 false (0.00) Sobol (0.04) false (0.00) false (0.07) 20 (0.04) BIPOP (0.00) mirrored (0.01) 10 (0.02) CSA (0.00) default (0.01) 0.396627 gaps/clusters
30 4 false (0.00) Sobol (0.04) false (0.00) false (0.06) 20 (0.03) BIPOP (0.00) mirrored (0.01) 10 (0.01) CSA (0.01) default (0.01) 0.382659 centre
30 5 false (0.06) Sobol (0.26) true (0.06) true (0.01) 5 (0.04) IPOP (0.00) mirrored (0.01) 5 (0.01) PSR (0.58) default (0.01) 0.992649 bounds
30 6 true (0.02) Sobol (0.14) false (0.03) false (0.38) 14 (0.00) BIPOP (0.00) mirrored (0.06) 7 (0.01) CSA (0.10) default (0.07) 0.583744 centre
30 7 true (0.00) Sobol (0.03) false (0.01) false (0.12) 14 (0.02) IPOP (0.01) pairwise (0.00) 5 (0.01) CSA (0.04) default (0.01) 0.447991 centre
30 8 true (0.01) Sobol (0.03) false (0.01) false (0.22) 5 (0.03) IPOP (0.01) mirrored (0.01) 5 (0.01) CSA (0.02) $\lambda$-decay (0.13) 0.409491 discretization
30 9 false (0.01) Sobol (0.01) true (0.16) true (0.06) 5 (0.02) IPOP (0.01) mirrored (0.00) 5 (0.01) CSA (0.00) equal (0.07) 0.409810 centre
30 10 true (0.00) Sobol (0.04) false (0.01) true (0.20) 10 (0.07) IPOP (0.00) pairwise (0.00) 10 (0.03) PSR (0.19) equal (0.09) 0.289573 gaps/clusters
30 11 true (0.01) Sobol (0.02) true (0.04) false (0.05) 14 (0.00) BIPOP (0.00) mirrored (0.03) 7 (0.01) PSR (0.01) default (0.01) 0.408763 centre
30 12 false (0.03) Sobol (0.08) true (0.51) true (0.04) 5 (0.17) BIPOP (0.06) mirrored (0.06) 5 (0.05) CSA (0.11) $\lambda$-decay (0.22) 0.505465 centre
30 13 true (0.03) Sobol (0.06) false (0.04) false (0.07) 14 (0.00) BIPOP (0.03) mirrored (0.04) 7 (0.03) PSR (0.03) default (0.05) 0.511117 centre
30 14 false (0.00) Gaussian (0.03) true (0.31) true (0.02) 5 (0.09) BIPOP (0.00) mirrored (0.01) 5 (0.01) CSA (0.03) $\lambda$-decay (0.14) 0.704631 centre
30 15 true (0.00) Sobol (0.02) false (0.01) false (0.03) 20 (0.04) IPOP (0.00) mirrored (0.01) 10 (0.02) PSR (0.00) default (0.01) 0.398682 centre
30 16 true (0.00) Sobol (0.03) false (0.01) false (0.10) 20 (0.06) IPOP (0.01) mirrored (0.04) 10 (0.02) CSA (0.01) default (0.03) 0.499925 centre
30 17 true (0.00) Sobol (0.08) false (0.02) false (0.20) 20 (0.10) BIPOP (0.00) mirrored (0.01) 10 (0.03) CSA (0.02) default (0.06) 0.635100 centre
30 18 true (0.00) Sobol (0.07) false (0.00) false (0.15) 20 (0.08) BIPOP (0.00) pairwise (0.02) 10 (0.03) CSA (0.03) default (0.02) 0.549818 centre
30 19 true (0.00) Gaussian (0.01) true (0.10) true (0.03) 5 (0.04) BIPOP (0.09) mirrored (0.00) 5 (0.00) CSA (0.10) equal (0.04) 0.548607 centre
30 20 true (0.00) Sobol (0.01) true (0.26) true (0.00) 5 (0.02) IPOP (0.00) mirrored (0.00) 5 (0.00) CSA (0.00) $\lambda$-decay (0.04) 0.481448 gaps/clusters
30 21 true (0.12) Sobol (0.16) true (0.28) true (0.18) 14 (0.00) IPOP (0.17) mirrored (0.18) 7 (0.12) CSA (0.15) $\lambda$-decay (0.14) 0.664094 centre
30 22 true (0.01) Gaussian (0.03) true (0.10) true (0.03) 5 (0.04) BIPOP (0.02) mirrored (0.03) 5 (0.01) CSA (0.04) $\lambda$-decay (0.06) 0.481261 centre
30 23 false (0.01) Sobol (0.02) true (0.00) false (0.09) 20 (0.05) IPOP (0.00) pairwise (0.01) 10 (0.01) PSR (0.09) default (0.01) 0.549599 centre
30 24 true (0.00) Sobol (0.01) false (0.00) false (0.03) 20 (0.02) IPOP (0.00) mirrored (0.00) 10 (0.01) PSR (0.01) default (0.01) 0.384299 centre
@@
The exact settings for each single-best, as well as the avg-best configuration, are shown in Table 5. This table highlights that, while the contribution trend of a module might be negative in Figure 2 or 3, it might still be beneficial in combination with other modules, or its impact might be sufficiently low when performance is high that changing its setting does not change the performance significantly. This can be observed for example with the ‘covariance’ parameter, which as mentioned before seems to have a positive contribution to performance when turned on. However, for 21 out of 50 settings, the best-performing configuration is an MA-ES instead of CMA-ES variant. Since Table 5 also shows the impact of each module choice relative to the average of the other choices, we can isolate the performance gained from enabling or disabling the covariance adaptation in more detail. We observe that in the cases where this adaptation is turned off, the normalized AOCC would be lowered by between $0.00$ and $0.03$ in 16 out of 21 cases, while the reverse change would lead to much larger deterioration in performance.
@@
Other patterns, such as the impact of mirrored sampling, are more consistent between the two types of analysis. Table 5 shows that all of the best configurations enable this module, with a slight preference for ‘mirrored’ over ‘mirrored pairwise’. The changes in performance from the single-module changes also highlight the low impact of restart strategies on unimodal problems. From looking only at the selected modules, we might infer that ‘BIPOP’ is needed to achieve optimal performance on F1, but since restarts are not triggered in this scenario, the performance of our selected configuration would not change if restart is changed to ‘IPOP’, or even turned off completely. The choice of elitist and local restart modules seem more stable regardless of dimensionality (the choice winning on a function for 5d is often also winning for that function in 30d). To use the covariance module and different population sizes however are not stable between the two dimensionalities.
@@ 4.4.1. Contribution of hyperparameters
Figure 4. Hyper-parameter contributions to the AOCC (shap values) per benchmark function for d=5 for modular DE. Options per module are sorted alphabetically (for categorical parameters) or numerically, refer to Table 2 for colour coding.
Figure 5. Hyper-parameter contributions per benchmark function for d=30 for modular DE. Options per module are sorted alphabetically (for categorical parameters) or numerically, refer to Table 2 for colour coding.
For the modular Differential Evolution, we have a total of 36 2883628836\,28836 288 configurations. We should note that, similar to the CMA-ES, the models from which we calculate the SHAP-values are intentionally overfit, resulting in $R^{2}$ scores between $0.9$ and $0.96$. Figures 4 and 5 show the resulting SHAP values for all functions, in dimensionality 5 and 30 respectively. For dimensionality 5, we observe strong effects from the population size parameter $\lambda$, with the default (smallest) value often contributing negatively to the anytime performance. For the unimodal problems, the population size $2\cdot d$ seems to be preferred, while the multimodal functions benefit from an increase to $10\cdot d$. For the adaptation method, we surprisingly find that turning this module off generally leads to good performance, while performance deteriorates when jDE is enabled.
@@
When looking at the discretized parameters F and CR, we observe some interesting effects. While for most functions, the impact changes are seemingly in trend with the size of the parameter, e.g., F becoming more positive the lower the value becomes, some 30-dimensional functions show completely different types of impact for CR. As an example, F7 shows that the highest CR value has both a very positive and a very negative impact on the anytime performance. This highlights an inherent limitation with this analysis method, in that we only focus on the single-module impacts. In this version of DE, is seems intuitive that several modules might interact with each other, such as the choice of crossover method and CR, which might explain the unusual SHAP plot for F7. Within IOHxplainer we can zoom in on these kinds of higher-order interactions by utilizing different explainability methods, we consider this outside the scope of this current paper.
@@ 4.4.2. Algorithm/configuration selection
Another interesting observation from Table 8 is that the highest population sizes are never chosen in the 30-dimensional setting. This is likely influenced by the combination of chosen budget (10 0001000010\,00010 000) and performance measure, where very high population sizes could lead to slower, but eventually more precise, convergence. This highlights that, as with any benchmarking setup, the design choices made should be carefully considered when drawing conclusions from IOHxplainer. Another interesting observation regarding population sizes is that the winning settings seem to be dimensionality-specific, as the winning configurations in 5d do not correspond to winning configurations in 30d (for the same function). In 5d smaller population sizes with no population size reduction (lpsr) is preferred on simpler functions while the opposite is true on more complex functions. In 30d, population size reduction seems less beneficial on most of the functions. The base-vector on the other hand is typically very stable over the different dimensionalities and should often be set to ‘target’.
@@
Table 7. Performance summary of modular DE. The gain avg-best is the average improvement of the average best configuration over all configurations. The gain single-best is the average improvement of the single-best configurations (per function) over all configurations. Boldface in the table indicates a significant improvement (avg-best versus all and single-best versus avg-best).
@@
Table 8. Hall of fame of modular DE: best configurations per function including average best (denoted by All in fid), with resulting values of AOCC and structural BIAS indicator. Denoted in brackets after each module setting is the average effect in AOCC score of choosing the respective option versus all other choices (keeping the rest of the settings the same). Colouring of module options is consistent with Table 2 and Figures 4,  5. Options marked in italic are consistently selected on this function for both dimensionalities.
@@
5 1 0.25 (0.25) 0.25 (0.29) none (0.12) bin (0.01) 8 (0.06) false (0.19) best (0.34) 1 (0.00) rand (0.02) true (0.55) 0.971233 gaps/clusters
5 2 0.25 (0.37) 0.25 (0.45) none (0.14) bin (0.01) 8 (0.13) false (0.38) best (0.31) 2 (0.00) rand (0.01) true (0.11) 0.940875 bounds
5 3 0.05 (0.40) 0.5 (0.59) none (0.13) exp (0.00) 8 (0.33) false (0.59) target (0.10) 2 (0.18) rand (0.08) true (0.06) 0.885237 -
@@
5 7 1.0 (0.35) 0.5 (0.44) none (0.08) bin (0.09) 50 (0.78) false (0.04) target (0.18) 1 (0.24) pbest (0.33) false (0.17) 0.909025 gaps/clusters
@@
5 9 0.75 (0.53) 0.5 (0.51) none (0.06) exp (0.01) 10 (0.43) true (0.01) target (0.29) 1 (0.09) pbest (0.16) true (0.66) 0.799625 -
5 10 1.0 (0.60) 0.25 (0.56) none (0.30) exp (0.04) 10 (0.49) true (0.01) target (0.76) 2 (0.25) none (0.74) true (0.39) 0.777871 -
5 11 1.0 (0.57) 0.25 (0.54) none (0.27) exp (0.01) 10 (0.42) true (0.01) target (0.76) 2 (0.17) none (0.69) true (0.26) 0.832987 -
@@
5 13 1.0 (0.49) 0.75 (0.55) none (0.20) bin (0.00) 50 (0.44) true (0.08) target (0.47) 1 (0.33) best (0.24) false (0.51) 0.714799 bounds
@@
5 16 1.0 (0.30) 0.25 (0.32) none (0.25) exp (0.08) 8 (0.22) true (0.29) best (0.33) 2 (0.10) rand (0.14) true (0.09) 0.466607 bounds
5 17 1.0 (0.33) 0.5 (0.34) none (0.16) exp (0.03) 50 (0.31) true (0.03) target (0.24) 1 (0.29) pbest (0.25) false (0.26) 0.713114 centre
5 18 1.0 (0.33) 0.5 (0.35) none (0.19) bin (0.01) 50 (0.31) true (0.02) target (0.25) 1 (0.30) pbest (0.29) false (0.25) 0.664711 centre
5 19 0.75 (0.04) 0.25 (0.05) jDE (0.06) exp (0.05) 10 (0.05) false (0.03) best (0.05) 1 (0.04) none (0.05) true (0.09) 0.259360 bounds
5 20 0.25 (0.25) 0.5 (0.26) shade (0.22) exp (0.13) 10 (0.37) true (0.16) rand (0.29) 1 (0.11) none (0.21) true (0.37) 0.678659 bounds
@@
5 23 0.75 (0.04) 0.25 (0.05) shade (0.05) exp (0.04) 50 (0.06) true (0.04) best (0.05) 2 (0.03) none (0.05) false (0.04) 0.236681 bounds
5 24 1.0 (0.03) 0.5 (0.05) none (0.02) bin (0.01) 50 (0.03) true (0.00) target (0.04) 1 (0.05) best (0.02) false (0.04) 0.123116 gaps/clusters
@@
30 2 0.25 (0.22) 0.75 (0.38) none (0.19) bin (0.19) 14 (0.38) false (0.32) target (0.32) 1 (0.24) pbest (0.13) false (0.14) 0.565194 -
30 3 0.05 (0.02) 0.5 (0.01) shade (0.01) exp (0.02) 14 (0.09) false (0.08) rand (0.03) 1 (0.02) none (0.03) true (0.00) 0.422156 -
30 4 0.05 (0.02) 0.5 (0.01) shade (0.01) exp (0.01) 14 (0.08) false (0.07) rand (0.03) 1 (0.02) none (0.02) true (0.00) 0.405671 -
30 5 0.25 (0.32) 1.75 (0.03) none (0.29) bin (0.05) 14 (0.23) false (0.11) best (0.02) 2 (0.14) best (0.02) true (0.03) 0.970231 bounds
@@
30 7 0.75 (0.03) 0.5 (0.06) none (0.01) bin (0.04) 60 (0.03) true (0.01) target (0.06) 1 (0.04) pbest (0.03) true (0.01) 0.416182 -
30 8 0.25 (0.06) 0.5 (0.12) none (0.01) bin (0.04) 14 (0.13) false (0.10) target (0.07) 2 (0.01) pbest (0.03) true (0.02) 0.375220 -
@@
30 11 1.0 (0.04) 0.25 (0.05) none (0.02) exp (0.00) 14 (0.01) true (0.03) target (0.03) 2 (0.01) pbest (0.02) true (0.01) 0.392524 centre
@@
30 16 0.75 (0.02) 0.25 (0.03) none (0.01) bin (0.02) 14 (0.02) false (0.02) best (0.03) 2 (0.01) none (0.00) true (0.01) 0.433995 bounds
30 17 0.5 (0.03) 0.5 (0.05) none (0.01) bin (0.05) 60 (0.02) true (0.01) target (0.04) 1 (0.04) pbest (0.03) false (0.00) 0.497815 -
30 18 0.75 (0.03) 0.5 (0.04) none (0.01) bin (0.04) 60 (0.01) true (0.01) target (0.04) 1 (0.03) pbest (0.03) false (0.01) 0.459689 centre
@@
30 20 0.25 (0.10) 0.25 (0.21) none (0.02) bin (0.02) 14 (0.14) false (0.08) best (0.02) 2 (0.01) rand (0.02) false (0.00) 0.472318 -
@@
30 23 1.0 (0.01) 0.25 (0.01) shade (0.02) exp (0.01) 14 (0.02) false (0.02) best (0.02) 2 (0.01) none (0.01) true (0.03) 0.481032 bounds